Errata for

R-SPACES ASSOCIATED WITH A HERMITIAN

SYMMETRIC PAIR

(Tsukuba J. Math. 10 (1986), 165–170)

Hiroyuki Tasaki

October 14, 1999 (revised, June 13, 2006 by Osami Yasukura)

p.167 Line 2 from bottom: “symmdtric pair” should read “symmetric pair”
p.168 Line 7: “M ∪ 𝔞” should read “M ∩ 𝔞”
p.169 Line 4: “υ [\frac{2\sqrt{−1}}{γ_{i}(H_{i})}]” should read “υ (\frac{2\sqrt{−1}}{γ_{i}(H_{i})}) H_{γ_{i}}”
p.169 Line 7: “ ∑_{{α∈Δ \atop α∈(v(H))=0}}” should read “ ∑_{{α∈Δ \atop α(v(H))=0}}”
p.169 Equation (5): “ ∑_{γ∉Γ_{H}}” should read “ ∑_{γ_{i}∉Γ_{H}}”